OpenStudy (anonymous):

Given the equation y - 3 = (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form x - 2y = -12 y = x y = x + 6 y = x + 9

OpenStudy (shamil98):

slope -intercept form is y = mx + b

OpenStudy (shamil98):

solve for y.

OpenStudy (anonymous):

I don't understand... Y would be 3?

OpenStudy (shamil98):

y - 3 = x + 6 solve for y. get y isolated on one side.

OpenStudy (shamil98):

3 is subtracted from y to get y back to its wholeself what would you do to do that?

OpenStudy (anonymous):